The number of ways of selecting 2 boys and 2 girls from a group of 4 boys and 5 girls

Answer by ewatrrr(24785)

 
Hi, Previously Posted
The number of ways of selecting 2 boys or 2 girls from a group of 4 boys and 5 girls
4C2*5C2 = 6*10 = 60 ways